OpenStudy (lukecrayonz):

derivative of lnsqrt(x+4)?

OpenStudy (anonymous):

remember a rule... derivative of ln (hathi) is 1/hathi . d(hathi) do that no n no. of terms

OpenStudy (anonymous):

\[\frac{ 1 }{ 2\sqrt{x+4} }\] Consider x+4 as a single term and differentiate...

OpenStudy (lukecrayonz):

@Salmon O.o can you explain that?

OpenStudy (lukecrayonz):

Been a few months since i've done any derivatives

OpenStudy (anonymous):

Cosider x+4=a so what is derivative of \[{\sqrt{a}}\] with respect to a?

OpenStudy (lukecrayonz):

1/2sqrt(x)

OpenStudy (lukecrayonz):

sorry a lol

OpenStudy (lukecrayonz):

not x

OpenStudy (anonymous):

k.good... Now what is the derivative of the term you considered as 'a' i.e x+4???

OpenStudy (anonymous):

with respect to x itself...

OpenStudy (lukecrayonz):

1

OpenStudy (lukecrayonz):

since x^1-0=0, anything to the power of 0 = 1

OpenStudy (anonymous):

Excellent, Now multiply both answer what will you get??

OpenStudy (lukecrayonz):

1/2sqrt(x)*1? same answer lol

OpenStudy (anonymous):

Now try solving (7x+9)^2

OpenStudy (lukecrayonz):

49x^2+126x+81 98x+126

OpenStudy (anonymous):

Good job dude..